# TIPO: Text to Image with Text Presampling for Prompt Optimization

### v3

@@ A.2 GBC10M
The GBC10M dataset [29] is a large-scale collection of 10 million images sourced from CC12M [13], annotated using the Graph-Based Captioning (GBC) approach. Each image is represented by a graph where nodes correspond to object regions, compositions, and relations, and edges define their hierarchical relationships. Annotations are generated automatically through a pipeline leveraging pretrained multimodal large language models (MLLM) and object detection tools. The GBC structure enhances traditional image captions by providing detailed descriptions and structural information. Data is provided in JSON lines format, including image URLs, bounding boxes, and captions.
@@ A.3 Coyo HD 11M
The Coyo HD 11M dataset [12] consists of 11.4 million high-resolution, high-concept-density images paired with 22.8 million synthetic captions generated from the Coyo-700M dataset. Images maintain a minimum of 512 pixels on the shortest edge to ensure high visual quality. Captions, generated with the LLaVA-Next-8B model [38] based on LLaMA 3 [3], undergo post-processing for conciseness and clarity.
@@
A.4 Stable Diffusion XL
Stable Diffusion XL (SDXL) [50] improves upon earlier models [56] with a more considerable UNet backbone and dual text encoders (CLIP ViT-L [51] and OpenCLIP ViT-bigG [31]), enhancing text conditioning. Supporting resolutions up to 1024×1024, SDXL accepts natural language prompts and tags, suitable for diverse image generation.
A.5 Illustrious
Illustrious is a series of fine-tuned Stable Diffusion XL models primarily trained on the Danbooru2023 dataset. In this study, we specifically employ the v3.5 version variant with v-parameterization [58], which is notable for its extensive incorporation of natural language prompts. The inclusion of both tag-based and natural language formats allows Illustrious to leverage a broad range of semantic knowledge for image generation.
@@ A.5 Illustrious
A.6 Stable Diffusion 3.5
Stable Diffusion 3.5 (SD-3.5) incorporates the MMDiT architecture [20] and the Rectified Flow formulation [39, 4, 36] for improved text-to-image generation. Utilizing triple text encoders (CLIP/ViT-L, OpenCLIP/ViT-G, T5-XXL [52]), SD-3.5 supports resolutions up to 1024×1024 and uses a 50/50 mix of original and CogVLM-generated captions. Figures confirm the capability to process both natural language prompts and tags.
@@
Appendix B Implementation Details
In this appendix, we provide all the necessary details including our dataset construction process, model configurations, inference pipeline and the model’s properties not mentioned in Section 4.2 and 4.3.
B.1 Training Data Construction
@@
B.2 Training Settings and Model Configurations
@@
Table 6: Training settings for TIPO models. The datasets include CoyoHD11M (Coyo), GBC10M (GBC), and Danbooru2023 (Dan). Stage 2 additionally incorporates Pixtral [1] to generate NL captions from Danboodu2023 dataset.
@@
TIPO employs a vocabulary derived from LLaMA2 [65] consisting of 32,000 tokens, with additional tokens (13 tokens) specifically designated for task and length control or placeholders. This extended vocabulary includes task identifiers and length modifiers to ensure flexibility across different prompt types:
@@
Training is performed using the AdamW optimizer [40], with a cosine annealing learning rate scheduler [41]. The optimizer parameters include $\beta_{1}=0.9$, $\beta_{2}=0.98$, and a weight decay of 0.01. Maximum learning rates are adjusted per model size, as outlined in Table 6.
@@
TIPO models are trained in multiple stages. Table 6 summarizes the configurations for pretraining and fine-tuning TIPO-100M, TIPO-200M, and TIPO-500M. Both pretraining and fine-tuning was conducted on datasets like Danbooru2023 [46], GBC10M [29], and CoyoHD11M [12].
@@
Figure 9: TIPO inference workflow, with solid arrows denoting the primary generation steps and dashed arrows indicating alternative generation paths within the same cycle. ¡TOKEN¿— represents special tokens, with all tokens detailed in Section B.2.
B.3 Inference Pipeline
The TIPO inference pipeline is designed to handle various input types and scenarios, combining different tasks to refine or expand both tag-based and natural language prompts. Figure 9 illustrates this comprehensive workflow.
Our framework processes tags and natural language inputs separately, allowing for specialized handling of each input type. This flexible pipeline allows TIPO to adapt to various input scenarios, whether the user provides tags, natural language descriptions, or both. By leveraging different task combinations, TIPO ensures that tag-based and natural language prompts are optimized, resulting in more detailed and effective input for text-to-image models.
B.4 Model Speed Comparison
@@
The evaluation was performed using llama.cpp [23], an efficient C++ implementation that provides optimized support for various hardware accelerators including CUDA, HIP, and Apple Metal.
@@
Appendix C Evaluation Statistics
@@
C.1 In-domain test regarding scenery tag
@@
C.2 In-domain prompt generation test
@@
Figures 11 and 12 display the box plots and KDE plots of aesthetic scores and AI corruption scores obtained from the In-domain prompt generation test detailed in Section E.2. While the box plots reveal subtle differences in performance between various methods, the AI corruption scores provide valuable insights. Specifically, these scores indicate that implementations supported by TIPO produce more stable output images than other methods.
C.3 Out-of-domain evaluation
@@
Figures 14 and 15 present similarity matrices for different prompt generation methods and their corresponding aesthetic outputs on SD3.5-Large  [21]. A matrix with predominantly lower similarity values (brighter appearance) indicates high diversity among generated images, while higher values (darker appearance) suggest consistent but less diverse outputs. Please refer to Table 3 in Section 5.
C.4 Ablation Test
@@
Appendix D TIPO example
@@
TIPO Format Template User Input: 1girl, ciloranko, maccha (mochancc), ningen mame, ask (askzy), solo, masterpiece, absurdres, newest, safe A girl sits in a cozy cafe, cradling a cup of coffee in her hand Formatted TIPO Input for Expand Tags: meta: absurdres rating: safe artist: ciloranko, maccha (mochancc), ningen mame, ask (askzy) quality: masterpiece, newest aspect_ratio: 1.0 target: <|short|> <|short_to_tag|> short: A girl sits in a cozy cafe, cradling a cup of coffee in her hand tag: 1girl, solo Formatted TIPO Output after Expand Tags and Expand Natural Prompt: meta: absurdres rating: safe artist: ciloranko, maccha (mochancc), ningen mame, ask (askzy) quality: masterpiece, newest aspect_ratio: 1.0 target: <|short|> <|tag_to_long|> tag: 1girl, solo, sitting, closed mouth, jewelry, long hair, looking at viewer, crossed legs, plant, table, couch, bracelet, cup, smile, teacup, indoors, blue eyes, blonde hair, holding long: A girl sits in a cozy cafe, cradling a cup of coffee in her hand. The cafe has large windows with green plants on the walls and a wooden table in front of her. Behind her is a staircase leading to another room. The overall atmosphere of the image is serene and inviting. Formatted Output for Text-to-Image: 1girl, ciloranko, maccha (mochancc), ningen mame, ask (askzy), solo, sitting, closed mouth, jewelry, long hair, looking at viewer, crossed legs, plant, table, couch, bracelet, cup, smile, teacup, indoors, blue eyes, blonde hair, holding, A girl sits in a cozy cafe, cradling a cup of coffee in her hand. Behind her is a staircase leading to another room. The cafe has large windows with green plants on the walls and a wooden table in front of her. The overall atmosphere of the image is serene and inviting. masterpiece, newest, absurdres, safe Refer to caption User Input Refer to caption TIPO Output
@@
Appendix E Image Examples
@@
E.1 In-domain test regard to scenery tag
@@
Figure 20 demonstrates the difference in output diversity between simple input and TIPO-enhanced input for the scenery tag. As observed, TIPO significantly expands the range of generated sceneries, better reflecting the variety present in the Danbooru2023 dataset [69]. The left column shows results from simple input (scenery tag only), while the right column illustrates the enhanced diversity achieved with TIPO-enhanced input.
E.2 In-domain prompt generation test
@@
Appendix F Human Preference
@@
Promptist vs. Tipo 30:101 0.229 $<0.0001$ Yes***
PromptExtend vs. Tipo 30:90 0.250 $<0.0001$ Yes***
@@
F.1 User Interface for Human Preference Evaluation
@@
F.2 Extended Human Evaluation.
@@
F.4 Human Preference ELO Method
@@
  • Method $i$ wins: assigned a score of $1$ for method $i$, and 00 for method $j$.
  • Method $j$ wins: assigned score $1$ for method $j$, and 00 for method $i$.
@@
Adjusted Win Rate=Win Rate+Tie $\text{Adjusted Win Rate}=\text{Win Rate}+\frac{\text{Tie Rate}}{2}$
ELO $\text{ELO Difference}=400\times\log_{10}\left(\frac{\text{Adjusted Win Rate}}{1-\text{Adjusted Win Rate}}\right)$
@@
ELO $\text{ELO Difference}=\begin{cases}-800,&\text{Adjusted Win Rate}\leq 0.001\\ +800,&\text{Adjusted Win Rate}\geq 0.999\\ \end{cases}$
@@
$\text{ELO}_{\text{method}_{i}}=\text{Base Rating}+\left(\text{Average ELO Difference for method }i-\text{Overall Mean ELO Difference}\right)$

### v4

@@ A.2 GBC10M
The GBC10M dataset (Hsieh et al., 2024) is a large-scale collection of 10 million images sourced from CC12M (Changpinyo et al., 2021), annotated using the Graph-Based Captioning (GBC) approach. Each image is represented by a graph where nodes correspond to object regions, compositions, and relations, and edges define their hierarchical relationships. Annotations are generated automatically through a pipeline leveraging pretrained multimodal large language models (MLLM) and object detection tools. The GBC structure enhances traditional image captions by providing detailed descriptions and structural information. Data is provided in JSON lines format, including image URLs, bounding boxes, and captions.
@@ A.3 Coyo HD 11M
The Coyo HD 11M dataset (CaptionEmporium, 2024) consists of 11.4 million high-resolution, high-concept-density images paired with 22.8 million synthetic captions generated from the Coyo-700M dataset. Images maintain a minimum of 512 pixels on the shortest edge to ensure high visual quality. Captions, generated with the LLaVA-Next-8B model (Liu et al., 2024a) based on LLaMA 3 (AI@Meta, 2024), undergo post-processing for conciseness and clarity.
@@
Appendix B Baselines/T2I models
B.1 Prompt-Optimization Baselines
GPT-4o-mini.
GPT-4o-mini(OpenAI, 2024) is a multimodal model introduced by OpenAI in July 2024 as a cost-efficient variant of GPT-4o. It supports a 128k-token context window and up to 16k output tokens, trained on text and vision data. In this paper, it serves as a zero-shot rewriting baseline for prompt refinement.
MagicPrompt.
MagicPrompt(daspartho, 2022) fine-tunes GPT-2 models on large-scale, community-collected prompts (e.g., from Lexica.art and the Stable Diffusion Prompts dataset). It learns to generate extended prompts with richer descriptive content, originally designed to improve prompt quality for Stable Diffusion.
Promptist.
Promptist(Hao et al., 2023) is a reinforcement learning framework for prompt optimization. Starting from a seed prompt, it generates refined prompts using supervised pre-training and reinforcement learning, guided by aesthetic and semantic alignment rewards. It consistently improves model-preferred prompt quality when paired with diffusion backbones.
Gemini-2.0-Flash-Image.
Gemini-2.0-Flash-Image [^4] https://developers.googleblog.com/experiment-with-gemini-20-flash-native-image-generation/ is a variant of Google’s Gemini 2.0 Flash family with native image generation support. It provides both prompt-refinement and text-to-image generation within a single model, supporting high-fidelity text rendering, compositional control, and iterative editing. Unlike other baselines, it functions as both optimizer and generator.
B.2 T2I Models
SDXL
Stable Diffusion XL (SDXL) (Podell et al., 2024) improves upon earlier models (Rombach et al., 2022b) with a more considerable UNet backbone and dual text encoders (CLIP ViT-L (Radford et al., 2021) and OpenCLIP ViT-bigG (Ilharco et al., 2021)), enhancing text conditioning. Supporting resolutions up to 1024×1024, SDXL accepts natural language prompts and tags, suitable for diverse image generation.
Illustrious
Illustrious is a series of fine-tuned Stable Diffusion XL models primarily trained on the Danbooru2023 dataset. In this study, we specifically employ the v3.5 version variant with v-parameterization (Salimans and Ho, 2022), which is notable for its extensive incorporation of natural language prompts. The inclusion of both tag-based and natural language formats allows Illustrious to leverage a broad range of semantic knowledge for image generation.
@@ B.2 T2I Models
Stable Diffusion 3.5
Stable Diffusion 3.5 (SD-3.5) incorporates the MMDiT architecture (Esser et al., 2024a) and the Rectified Flow formulation (Liu et al., 2022; Albergo and Vanden-Eijnden, 2022; Lipman et al., 2022) for improved text-to-image generation. Utilizing triple text encoders (CLIP/ViT-L, OpenCLIP/ViT-G, T5-XXL (Raffel et al., 2020b)), SD-3.5 supports resolutions up to 1024×1024 and uses a 50/50 mix of original and CogVLM-generated captions. Figures confirm the capability to process both natural language prompts and tags.
@@
FLUX.1-dev
FLUX.1-dev is an open-weights text-to-image model released by Black Forest Labs as a guidance-distilled variant of their FLUX.1 family, targeting research and non-commercial use (black-forest-labs, 2024). Architecturally, FLUX adopts a transformer-based rectified-flow formulation (Lipman et al., 2022) with a hybrid stack of multimodal/parallel diffusion transformer blocks and rotary position encodings, scaled to $\sim$12B parameters, emphasizing prompt adherence, typography, and aspect-ratio flexibility (black-forest-labs, 2024). We use the FLUX.1-dev checkpoint for T2I without additional refiners; it accepts both natural-language prompts and tag-like inputs and supports resolutions in the 0.10.12.02.0 MP range. [^10] https://huggingface.co/black-forest-labs/FLUX.1-dev
OmniGen2
OmniGen2 is a unified multimodal generator that decouples autoregressive text modeling from diffusion-based image generation via two distinct pathways with unshared parameters (Wu et al., 2025). The diffusion side conditions on hidden states from the MLLM while exclusively feeding VAE features into the diffusion decoder to preserve low-level fidelity; a 3D rotary scheme (Omni-RoPE) disentangles sequence ID and 2D spatial coordinates to stabilize editing and in-context generation (Wu et al., 2025). Beyond standard T2I, OmniGen2 natively supports image editing and subject-driven in-context generation, and introduces a reflection mechanism/dataset to iteratively refine outputs.
Lumina-Image 2.0
[…]
HiDream-I1
HiDream-I1 is a 17B-parameter image foundation model based on a sparse Diffusion Transformer with dynamic Mixture-of-Experts (MoE) (Cai et al., 2025). It employs a dual-stream (text/image) sparse DiT for separate encoding followed by a single-stream sparse DiT to fuse modalities efficiently; hybrid text encoders (e.g., CLIP-L/14, CLIP-G/14, T5-XXL) and an LLM aggregator provide robust conditioning (Cai et al., 2025). The suite includes I1-Full (50+50{+} steps), I1-Dev (guidance-distilled, 28 steps), and I1-Fast (14 steps), with a GAN-powered diffusion distillation to retain sharpness at low step counts.
Appendix C TIPO Implementation Details
In this appendix, we provide all the necessary details including our dataset construction process, model configurations, inference pipeline, and the model’s properties not mentioned in Section 4.2 and 4.3.
C.1 TIPO Training Data Construction
@@
C.2 TIPO Training Settings and Model Configurations
@@
Table 8: Training settings for TIPO models. The datasets include CoyoHD11M (Coyo), GBC10M (GBC), and Danbooru2023 (Dan). Stage 2 additionally incorporates Pixtral (Agrawal et al., 2024) to generate NL captions from Danbooru2023 dataset.
@@
TIPO employs a vocabulary derived from LLaMA2 (Touvron et al., 2023b) consisting of 32,000 tokens, with additional tokens (13 tokens) specifically designated for task and length control or placeholders. This extended vocabulary includes task identifiers and length modifiers to ensure flexibility across different prompt types:
@@
Training is performed using the AdamW optimizer (Loshchilov, 2017), with a cosine annealing learning rate scheduler (Loshchilov and Hutter, 2017). The optimizer parameters include $\beta_{1}=0.9$, $\beta_{2}=0.98$, and a weight decay of 0.01. Maximum learning rates are adjusted per model size, as outlined in Table 8.
@@
TIPO models are trained in multiple stages. Table 8 summarizes the configurations for pretraining and fine-tuning TIPO-100M, TIPO-200M, and TIPO-500M. Both pretraining and fine-tuning was conducted on datasets like Danbooru2023 (nyanko202, 2023), GBC10M (Hsieh et al., 2024), and CoyoHD11M (CaptionEmporium, 2024).
@@
Figure 6: TIPO inference workflow, with solid arrows denoting the primary generation steps and dashed arrows indicating alternative generation paths within the same cycle. ¡TOKEN¿— represents special tokens, with all tokens detailed in Section C.2.
C.3 TIPO Inference Settings
Sampling Strategy
We employ a hybrid stochastic decoding strategy combining nucleus sampling (top-$p$ = 0.95) and top-$k$ = 60 filtering. This follows standard practice in open-ended text generation, as adopted in the official Hugging Face generation examples  [^11] Hugging Face. “Usage — transformers 2.11.0 documentation.” Example of text generation with XLNet uses top-$p$ = 0.95 and top-$k$ = 60. https://huggingface.co/transformers/v2.11.0/usage.html. This hybrid approach maintains diversity while preserving coherence, preventing both overly deterministic and excessively noisy generations.
Inference Pipeline
The TIPO inference pipeline is designed to handle various input types and scenarios, combining different tasks to refine or expand both tag-based and natural language prompts. Figure 6 illustrates this comprehensive workflow. Our framework processes tags and natural language inputs separately, allowing for specialized handling of each input type. This flexible pipeline allows TIPO to adapt to various input scenarios, whether the user provides tags, natural language descriptions, or both. By leveraging different task combinations, TIPO ensures that tag-based and natural language prompts are optimized, resulting in more detailed and effective input for text-to-image models.
C.4 Impact of Model Size on Performance
To analyze the impact of model scales on prompt-optimization performance, we compare TIPO-200M and TIPO-500M using a 1,000-image subsample from the COYO and GBC datasets. Results are shown in Table 9.
Table 9: Prompt optimization performance of TIPO-200M and TIPO-500M on a 1k subsample.
Metric Task TIPO-200M TIPO-500M
FDD (↓) NL-short 0.1529 0.1356
NL-long 0.1650 0.1398
Aesthetic (↑) NL-short 5.8531 $\pm$ 0.7501 5.8943 $\pm$ 0.7064
NL-long 5.8364 $\pm$ 0.7501 5.9030 $\pm$ 0.7015
AI Corrupt (↓) NL-short 0.2870 $\pm$ 0.4167 0.2891 $\pm$ 0.4189
NL-long 0.2870 $\pm$ 0.4150 0.2862 $\pm$ 0.4151
Overall, TIPO-500M shows consistent gains in FDD and Aesthetic scores, while performance on AI Corrupt remains comparable. However, the larger 500M variant entails substantially higher computational cost without delivering proportionally greater benefits, which limits its practicality for community use; hence, all main experiments are conducted with TIPO-200M.
C.5 Impact of Model Size on Inference Speed
@@
The evaluation was performed using llama.cpp (Gerganov, 2023), an efficient C++ implementation that provides optimized support for various hardware accelerators, including CUDA, HIP, and Apple Metal.
@@
Appendix D Evaluation Statistics
@@
D.1 In-domain test regarding scenery tag
@@
D.2 In-domain prompt generation test
@@
Figures 8 and 9 display the box plots and KDE plots of aesthetic scores and AI corruption scores obtained from the In-domain prompt generation test detailed in Section F.2. While the box plots reveal subtle differences in performance between various methods, the AI corruption scores provide valuable insights. Specifically, these scores indicate that implementations supported by TIPO produce more stable output images than other methods.
D.3 Out-of-domain evaluation
@@
Figures 11 and 12 present similarity matrices for different prompt generation methods and their corresponding aesthetic outputs on SD3.5-Large  (Esser et al., 2024b). A matrix with predominantly lower similarity values (brighter appearance) indicates high diversity among generated images, while higher values (darker appearance) suggest consistent but less diverse outputs. Please refer to Table 2 in Section 5.
D.4 Ablation Test
@@
Appendix E TIPO example
@@
TIPO Format Template User Input: 1girl, ciloranko, maccha (mochancc), ningen mame, ask (askzy), solo, masterpiece, absurdres, newest, safe A girl sits in a cozy cafe, cradling a cup of coffee in her hand Formatted TIPO Input for Expand Tags: meta: absurdres rating: safe style: anime-style illustration, digital art, character design, fantasy concept art quality: masterpiece, newest aspect_ratio: 1.0 target: <|short|> <|short_to_tag|> short: A girl sits in a cozy cafe, cradling a cup of coffee in her hand tag: 1girl, solo Formatted TIPO Output after Expand Tags and Expand Natural Prompt: meta: absurdres rating: safe style: anime-style illustration, digital art, character design, fantasy concept art quality: masterpiece, newest aspect_ratio: 1.0 target: <|short|> <|tag_to_long|> tag: 1girl, solo, sitting, closed mouth, jewelry, long hair, looking at viewer, crossed legs, plant, table, couch, bracelet, cup, smile, teacup, indoors, blue eyes, blonde hair, holding long: A girl sits in a cozy cafe, cradling a cup of coffee in her hand. The cafe has large windows with green plants on the walls and a wooden table in front of her. Behind her is a staircase leading to another room. The overall atmosphere of the image is serene and inviting. Formatted Output for Text-to-Image: 1girl, ciloranko, maccha (mochancc), ningen mame, ask (askzy), solo, sitting, closed mouth, jewelry, long hair, looking at viewer, crossed legs, plant, table, couch, bracelet, cup, smile, teacup, indoors, blue eyes, blonde hair, holding, A girl sits in a cozy cafe, cradling a cup of coffee in her hand. Behind her is a staircase leading to another room. The cafe has large windows with green plants on the walls and a wooden table in front of her. The overall atmosphere of the image is serene and inviting. masterpiece, newest, absurdres, safe Refer to caption User Input Refer to caption TIPO Output
@@
Appendix F Image Examples
@@
F.1 In-domain test regard to scenery tag
@@
Figure 17 demonstrates the difference in output diversity between simple input and TIPO-enhanced input for the scenery tag. As observed, TIPO significantly expands the range of generated sceneries, better reflecting the variety present in the Danbooru2023 dataset (Yeh, 2024a). The left column shows results from simple input (scenery tag only), while the right column illustrates the enhanced diversity achieved with TIPO-enhanced input.
F.2 In-domain prompt generation test
@@
Appendix G Human Preference
@@
Promptist vs. Tipo 30:101 0.229 <0.0001<0.0001 Yes***
PromptExtend vs. Tipo 30:90 0.250 <0.0001<0.0001 Yes***
@@
G.1 User Interface for Human Preference Evaluation
@@
G.2 Extended Human Evaluation.
@@
G.4 Human Preference ELO Method
@@
  • Method $i$ wins: assigned a score of $1$ for method $i$, and 0 for method $j$.
  • Method $j$ wins: assigned score $1$ for method $j$, and 0 for method $i$.
@@
$\text{Adjusted Win Rate}=\text{Win Rate}+\frac{\text{Tie Rate}}{2}$
$\text{ELO Difference}=400\times\log_{10}\left(\frac{\text{Adjusted Win Rate}}{1-\text{Adjusted Win Rate}}\right)$
@@
[…]
@@
[…]